Dany contributes property to a new partnership for a 50 percent interest in the partnership. The property has an adjusted basis in Dany's hands of $30,000. The property was also subject to an $80,000 mortgage that was assumed by the partnership. What is Dany's basis in his partnership interest?
Answer #1

Dany has 50 percent interest in the partnership.

His net liablity relief = 80000 * 50 % = $ 40000

Dany's basis = 30000 - 40000 = 0